OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. I need HELP

OpenStudy (anonymous):

A. 6 B. -6 C.All real numbers D. No solution

OpenStudy (horsegirl27):

I think I can help

OpenStudy (horsegirl27):

What math is this?

OpenStudy (anonymous):

@iGreen or @mathmath333 could you help me plz?

OpenStudy (horsegirl27):

I can help you

OpenStudy (anonymous):

So you need to start out simplifying. Can you disribute the '-3' in front of the brackets?

OpenStudy (chosenmatt):

the equation is false so that means thiers no solution i think

OpenStudy (horsegirl27):

So first do distributive property. What do you get? I want you to try it

OpenStudy (anonymous):

ok thx guys :)

OpenStudy (horsegirl27):

You want to distribute -3

OpenStudy (horsegirl27):

On both sides

OpenStudy (chosenmatt):

i think it's no solution but maybe @horsegirl27 can confirm my answer

OpenStudy (horsegirl27):

After distributing, combine like terms and see what u get

OpenStudy (horsegirl27):

@chosenmatt yes lets work it out and see

OpenStudy (anonymous):

ok but i dont get what all real numbers means, and i know its not 6

OpenStudy (horsegirl27):

I want @Calvin3 to work it out and learn how to do it first

OpenStudy (horsegirl27):

Real numbers are all numbers on the number line, including rational and irrational numbers

OpenStudy (anonymous):

@Nerd94 can you do this?

OpenStudy (horsegirl27):

Now lets not jump ahead to the answer. Slow down and work out your problem to quickly find the answer. Stop guessing

OpenStudy (anonymous):

wouldnt you just plug in one of the answers too see if it works? and thx @horsegirl27 for helping me with the real numbers thing

OpenStudy (horsegirl27):

-3(x+3) = -3(x+1)-5 -3x -9 = -3x -3 -5 you should see your answer in there already

OpenStudy (anonymous):

so would the answer be no solution because i just did the work and -6 isn't the answer.

OpenStudy (horsegirl27):

what is the next step? I want you to do it. It's not hard and your not supposed to tell people just the answer.

OpenStudy (horsegirl27):

Yes

OpenStudy (horsegirl27):

@chosenmatt was right but I wanted you to try it so you would understand how to do it

OpenStudy (chosenmatt):

thanks @horsegirl27

OpenStudy (horsegirl27):

At least I got D, no solution

OpenStudy (chosenmatt):

i gave u a medal

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ -3(x + 3) = -3(x + 1) - 5 \hspace{.33em}\\~\\ -3x - 9 = -3x-3 - 5 \hspace{.33em}\\~\\ \cancel {-3x} - 9 = \cancel{ -3x} -3 - 5 \hspace{.33em}\\~\\ -9\neq -8 }\end{align}\) see if it makes sense

OpenStudy (horsegirl27):

Thanks @mathmath333 now I don't have to type it!

OpenStudy (anonymous):

i got No solution too D.

OpenStudy (horsegirl27):

Good. I'm glad you tried it on your own to check.

OpenStudy (anonymous):

After simplifying, you'll reach -1 = 0, which is just false. Whatever the value of x is, the equation will always be false and therefore there are no solutions.

OpenStudy (anonymous):

so do you agree its D.?

OpenStudy (horsegirl27):

Yes

OpenStudy (anonymous):

well thx guys for all your help, all you guys where kind to help me out!!!:)

OpenStudy (horsegirl27):

Your welcome